Tell whether if the following piecewise function is a continuous at a given point or not. (SHOW THE SOLUTION).



1. at x = 4



x + 1 if x < 4


(x - 4)² + 3 if x ≥ 4




1
Expert's answer
2022-06-08T13:38:49-0400

1.


"\lim\limits_{x\to4^-}f(x)=\lim\limits_{x\to4^-}(x+1)=4+1=5""\lim\limits_{x\to4^+}f(x)=\lim\limits_{x\to4^+}((x-4)^2+3)=(4-4)^2+3=3""\lim\limits_{x\to4^-}f(x)=5\not=3=\lim\limits_{x\to4^+}f(x)""\lim\limits_{x\to4}f(x)=\text{does not exist}"

The function "f(x)" is not continuous at "x=4."

The function "f(x)" has a jump discontinuity at "x=4."